OpenStudy (anonymous):

Mr. Matthews is conducting job interviews. He has 5 candidates for a teaching job and must choose 3 of them to go on to the second round of interviews. How many possible combinations of individuals exist?

OpenStudy (anonymous):

5!/(3!*2!) which equals (5*4*3*2*1)/(3*2*1*2*1)= 5*2=10. I think this is right, but i might have forgotten how to do these.

OpenStudy (anonymous):

Yes, that way worked aswell..but I did the same for the 1st part then did the other half of the formula & divided & got the same answer. But thanks.

OpenStudy (anonymous):

The way i did it was 5 choose 3, or on the calculator 5 nCr 3. which just makes these type of questions faster.

OpenStudy (anonymous):

I see what you're saying. Thankyou.